OpenStudy (anonymous):

square root of x= √2x-33

OpenStudy (mathmate):

Do you mean to solve for x in: \[x=\sqrt{2x-33}\]

OpenStudy (anonymous):

\[x^2=2x-33\] \[x^2-2x+33=0\]

OpenStudy (anonymous):

do you mean \[\sqrt{x}=\sqrt{2x-33}\]?

OpenStudy (anonymous):

I mean Taking the Square root of x²=2x+33 which I put into x= √2x-33 and now I'm stuck.

OpenStudy (binary3i):

\[[2pmsqrt{-128}]/2 complex roots\]

OpenStudy (anonymous):

x^2-2x+33=0 You will need to use the quadratic formula to solve this

OpenStudy (mathmate):

x^2-2x+33=0 gives complex roots. Do you work with complex roots, Mandy?

OpenStudy (anonymous):

\[x^{2}-2x+33=(x-1)^{2} +31>=31 >0 \] so it dont have real roots

OpenStudy (mathmate):

"I mean Taking the Square root of x²=2x+33 which I put into x= √2x-33" We don't really understand by your statement above. You can help to clarify by putting parentheses. For example, "square root of x^2" or "square root of (x^2=2x+33)"? I do not understand either interpretation.

OpenStudy (earthcitizen):

\[x ^{2}=2x-33 => x=\sqrt{(2x-33)}=>x=(2x-33)^{1/2}\]

OpenStudy (earthcitizen):

to solve for the roots of x \[x=\sqrt{2x-33} => x ^{2}-2x+33=0 (Zed )\] Using,\[x=(-b \pm \sqrt{b ^{2}-4ac})/2a\] \[a=1, b=-2, c=33\] x=\[(-(-2)\pm \sqrt{(-2)^2-4(1)*(33)})/2(1)\] x=\[(2\pm \sqrt{-128})/2\]

OpenStudy (earthcitizen):

\[x=(2+j8\sqrt{2}), (2-j8\sqrt{2})/2 => x=1+j4\sqrt{2}, 1-j4\sqrt{2}\]